OpenStudy (oswaldmurphy):

Given the function f(x) = 5x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. (4 points) Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. (6 points)

OpenStudy (oswaldmurphy):

@TheSmartOne @jim_thompson5910

OpenStudy (oswaldmurphy):

Would A be 7.5?

OpenStudy (oswaldmurphy):

@.Sam. @sammixboo @agent0smith

jimthompson5910 (jim_thompson5910):

how are you getting 7.5?

OpenStudy (oswaldmurphy):

0 times 5 is 0 and 1 times 5 is 5. 2.5. 2 times 5 is 10 and 3 times 5 is 15. So uh 5 then..?

jimthompson5910 (jim_thompson5910):

f(x) = 5x f(0) = 5*0 f(0) = 0 f(x) = 5x f(1) = 5*1 f(1) = 5 So we have the two points (0,0) and (1,5)

jimthompson5910 (jim_thompson5910):

find the slope through (0,0) and (1,5) to find the average rate of change for section A

OpenStudy (oswaldmurphy):

slope is 5. Rate of change is 5?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what about the slope of section B?

jimthompson5910 (jim_thompson5910):

wait, is the function f(x) = 5x OR is it f(x) = 5^x ??

OpenStudy (oswaldmurphy):

(2, 10) (3, 15) so 5?

OpenStudy (oswaldmurphy):

5^x

jimthompson5910 (jim_thompson5910):

ok be careful about inserting exponents

jimthompson5910 (jim_thompson5910):

redo section A and B again with this new function

OpenStudy (oswaldmurphy):

(0, 5) (1, 25) ?

jimthompson5910 (jim_thompson5910):

section B is from x = 2 to x = 3

OpenStudy (oswaldmurphy):

Is section a correct?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (oswaldmurphy):

Wait what does 5^1 =?

jimthompson5910 (jim_thompson5910):

actually, no, sorry

jimthompson5910 (jim_thompson5910):

f(x) = 5^x f(0) = 5^0 f(0) = 1

jimthompson5910 (jim_thompson5910):

5^0 is not 0

OpenStudy (oswaldmurphy):

Its 5.

OpenStudy (oswaldmurphy):

oh nvm..

OpenStudy (oswaldmurphy):

(0, 1) and (1, 5) for a?

jimthompson5910 (jim_thompson5910):

as stated earlier, if you plug in x = 0 into f(x) = 5^x, you get 1 as a result

jimthompson5910 (jim_thompson5910):

oh nvm I see what you mean now

jimthompson5910 (jim_thompson5910):

yes section A is from (0,1) to (1,5). Compute the slope of this

OpenStudy (oswaldmurphy):

4.

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

how about section B?

OpenStudy (oswaldmurphy):

(2, 25) (3, 625) 600 is slope for that

jimthompson5910 (jim_thompson5910):

5^3 isn't 625

OpenStudy (oswaldmurphy):

oh oops 125, so slope is 100

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

`How many times greater is the average rate of change of Section B than Section A?` fill in the blank B = _______*A where B refers to "slope of section B" and A refers to "slope of section A"

OpenStudy (oswaldmurphy):

A: For section A its 4, and section B its 100. B: 25 times. Its much bigger because x = 2 and x = 3 are bigger than x = 0 and x = 1.

OpenStudy (oswaldmurphy):

Is that correct?

jimthompson5910 (jim_thompson5910):

I agree with the first part of your answer in part B, but you should be more specific in the second part of part B

jimthompson5910 (jim_thompson5910):

`Its much bigger because x = 2 and x = 3 are bigger than x = 0 and x = 1.` is too vague in my opinion

OpenStudy (oswaldmurphy):

Is that fine for an assignment when you are 13? XD

jimthompson5910 (jim_thompson5910):

I guess it depends on your teacher, but I'd include the term "exponential function" in there somewhere

OpenStudy (oswaldmurphy):

I dont know what that means.

jimthompson5910 (jim_thompson5910):

Exponential growth functions have a property that the average rate of change increases over time

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):